# On a $T_{1}$ Transport inequality for the adapted Wasserstein distance

### v1

@@ Abstract.
The $L^{1}$ transport-entropy inequality (or $T_{1}$ inequality), which bounds the 111-Wasserstein distance in terms of the relative entropy, is known to characterize Gaussian concentration. To extend the $T_{1}$ inequality to laws of discrete-time processes while preserving their temporal structure, we investigate the adapted $T_{1}$ inequality which relates the 111-adapted Wasserstein distance to the relative entropy. Building on the Bolley–Villani inequality, we establish the adapted $T_{1}$ inequality under the same moment assumption as the classical $T_{1}$ inequality.
@@ 1. Introduction
Here, $H\left(\cdot\,|\,\mu\right)$ denotes the relative entropy with respect to $\mu$, defined as
[…]
There are well-established connections between $T_{p}$ inequalities and concentration of measure. In particular, when $p=1$, it is shown in [8, 10] that the following three statements are equivalent:
@@
    $\mu\in\mathcal{P}_{1}(\mathcal{X})$ satisfies the $T_{1}$ inequality;
@@
    $\mu$ exhibits Gaussian concentration; see [15, Theorem 22.10];
@@
    There exist $x_{0}\in\mathcal{X}$ and $\alpha>0$ such that
@@
The constant $C$ appearing in the $T_{1}$ inequality (2) can be explicitly expressed in terms of the exponential moment of $\mu$ in (4). More precisely, Bolley–Villani [9] established the following bound:
(5) $\mathcal{W}_{1}(\mu,\nu)\leq\frac{1}{\sqrt{\alpha}}\left(1+\log\int_{\mathcal{X}}e^{\alpha d(x,x_{0})^{2}}\mu(dx)\right)^{1/2}\sqrt{2H\left(\nu\,|\,\mu\right)}$
for all $\mu,\nu\in\mathcal{P}(\mathcal{X})$, $x_{0}\in\mathcal{X}$ and $\alpha>0$. This is a special case of the following bound on the weighted total variation, obtained by choosing $\varphi(x)=\sqrt{\alpha}d(x,x_{0})$ below.
@@
Our goal is to extend the Bolley–Villani bound (5) to the adapted Wasserstein distance when $\mu$ and $\nu$ are laws of discrete-time processes. This will follow from the adapted version of Theorem 1 which we present in Theorem 6.
@@ 1.1. Main results
While the topology induced by $\mathcal{A}\mathcal{W}_{p}$ is well understood (see [5]), we emphasize that it is the smallest topology under which optimal stopping problems have continuous value functions.
Now, we define the adapted weighted total variation and present our main result, Theorem 6.
@@
where the infimum is attained when $\pi(dx,dy)=(\operatorname{id},\operatorname{id})_{\#}(\mu\wedge\nu)$ on $\{x=y\}$. For example, see [13, Section 2] or [1, Lemma 3.1]. In this context, $\operatorname{ATV}_{\varphi}$ can be viewed as a natural extension of $\operatorname{TV}_{\varphi}$. The study of $\operatorname{ATV}_{\varphi}$ is not new and related works will be discussed in the next section.
@@
By properly choosing the weight function $\varphi$, we obtain the transport-entropy inequality for $\mathcal{A}\mathcal{W}_{1}$.
@@
Let $\mu\in\mathcal{P}_{1}(\mathcal{X})$. Then the following two properties are equivalent:
@@
    $\mu$ satisfies the adapted $T_{1}$ inequality, i.e, there exists some constant $C>0$ such that
If these properties hold, we have $\int_{\mathcal{X}}e^{\alpha d(x,x_{0})^{2}}\mu(dx)<\infty$ for some $x_{0}\in\mathcal{X}$ and $\alpha>0$ and
(13) $\mathcal{A}\mathcal{W}_{1}(\mu,\nu)\leq\frac{2\sqrt{T}+1}{\sqrt{\alpha}}\left(1+\log\int_{\mathcal{X}}e^{\alpha d(x,x_{0})^{2}}\mu(dx)\right)^{1/2}\sqrt{2H\left(\nu\,|\,\mu\right)}\text{ for all }\nu\in\mathcal{P}(\mathcal{X}).$
[…]
@@ 1.2. Related works
Let us recall that if the reference measure $\mu$ has independent marginals, i.e, $\mu=\otimes_{t=1}^{T}\gamma_{t}$ and each $\gamma_{t}$ satisfies the $T_{1}$ inequality (2) with a constant $C>0$, then the tensorization argument yields
@@
Our result is based on the Bolley–Villani bound (Theorem 1) and does not extend the tensorization technique beyond the independent case. We aim to investigate its generalization in the future research.
@@
(21) $\operatorname{TV}(\mu,\nu):=2\inf_{\operatorname{Cpl}(\mu,\nu)}\pi(x\neq y)\leq\sqrt{2H\left(\nu\,|\,\mu\right)}.$
@@
holds with $C=2^{T}-1$. The optimal constant $C=2T-1$ was later obtained by Acciaio et al. [1]. In combination with the classical Pinsker inequality, this implies
@@
Acciaio et al. [1] further showed that if $\mu\in\mathcal{P}((\mathbb{R}^{d})^{T})$ satisfy
(27) $\mu\text{-ess sup}\frac{1}{1+\sum_{s=1}^{t}\left|x_{t}\right|}\int\left|x_{t+1}\right|\mu^{x_{1},\ldots,x_{t}}(dx_{t+1})<\infty\text{ for all }t\in\{1,2,\ldots,T-1\},$
@@
for some constant $C>0$ that rises from the assumption (27). This estimate, when combined with the Bolley–Villani bound (Theorem 1), naturally leads to the following transport inequality:
@@ 2. Proofs
where $\eta_{1}:=\mu_{1}\wedge\nu_{1}$. For each $t\in\{1,2,\ldots,T-1\}$ and $x_{1:t},y_{1:t}\in\mathcal{X}_{1:t}$, we define $\gamma^{x_{1:t},y_{1:t}}\in\operatorname{Cpl}(\mu^{x_{1:t}},\nu^{y_{1:t}})$ via
@@
The second inequality follows from the minimality of $\mu_{1:j-1}\wedge\nu_{1:j-1}$. The proof of the last case $j=T$ is essentially the same. Indeed,
@@
(61) $\leq\int_{\mathcal{X}}\varphi(x)\left|\mu^{x_{1:T-1}}-\nu^{x_{1:T-1}}\right|(dx_{T})\mu_{1:T-1}\wedge\nu_{1:T}(dx_{1:T-1})$
@@
(78) $=\left(\int_{\mathcal{X}_{1:j-1}}f(x_{1:j-1})^{1/2}g(x_{1:j-1})^{1/2}\mu_{1:j-1}\wedge\nu_{1:j-1}(dx_{1:j-1})\right)^{2}$
@@
[…]

### v2

@@ Abstract.
The $L^{1}$ transport-entropy inequality (or $T_{1}$ inequality), which bounds the $1$-Wasserstein distance in terms of the relative entropy, is known to characterize Gaussian concentration. To extend the $T_{1}$ inequality to laws of discrete-time processes while preserving their temporal structure, we investigate the adapted $T_{1}$ inequality which relates the $1$-adapted Wasserstein distance to the relative entropy. Building on the Bolley–Villani inequality, we establish the adapted $T_{1}$ inequality under the same moment assumption as the classical $T_{1}$ inequality.
@@ 1. Introduction
If (2) holds with constant $C$, then $\mu$ is said to satisfy the $T_{p}(C)$ inequality. Here, $H\left(\cdot\,|\,\mu\right)$ denotes the relative entropy with respect to $\mu$, defined as
(3) $H\left(\nu\,|\,\mu\right):=\begin{cases}\int_{\mathcal{X}}\log\left(\frac{d\nu}{d\mu}\right)d\nu&\text{ if }\nu\ll\mu,\\ +\infty&\text{ otherwise.}\end{cases}$
There are well-established connections between $T_{p}$ inequalities and concentration of measure. In particular, when $p=1$, it is shown in [8, 10] that the $T_{1}$ inequality is equivalent to Gaussian concentration.
Theorem 1 (Theorem $22.10$ in [15]).
Let $\mu\in\mathcal{P}_{1}(\mathcal{X})$ and $C>0$. Each of the following statements implies the next one:
@@
    $\mu$ satisfies the $T_{1}(C)$ inequality.
@@
    $\mu$ exhibits Gaussian concentration: for any $\varepsilon>0$ and $1$-Lipschitz $f:\mathcal{X}\to\mathbb{R}$ such that $f\in L^{1}(\mu)$,
    (4) $\mu(\{x\in\mathcal{X}:f(x)\geq\int_{\mathcal{X}}fd\mu+\varepsilon\})\leq e^{-\frac{\varepsilon^{2}}{2C^{2}}}.$
@@
    $\mu$ has a finite quadratic exponential moment: for any $x_{0}\in\mathcal{X}$ and $\alpha\in(0,\frac{1}{2C^{2}})$,
@@
  4. (d)
    For any $x_{0}\in\mathcal{X}$ and $\alpha\in(0,\frac{1}{2C^{2}})$, $\mu$ satisfies the $T_{1}(C^{\prime})$ inequality where
    (6) $C^{\prime}=\frac{1}{\sqrt{\alpha}}\left(1+\log\int_{\mathcal{X}}e^{\alpha d(x,x_{0})^{2}}\mu(dx)\right)^{1/2}<\infty.$
The last implication from $(c)$ to $(d)$ is a special case of the following bound on the weighted total variation, obtained by choosing $\varphi(x)=\sqrt{\alpha}d(x,x_{0})$ below.
@@
Here, $\left|\mu-\nu\right|$ denotes the total variation measure of $\mu-\nu$ defined via
(8) $\left|\mu-\nu\right|(A)=\int_{A}\left|\frac{d\mu}{d(\mu+\nu)}-\frac{d\nu}{d(\mu+\nu)}\right|d(\mu+\nu)$
for all measurable $A$.
Our goal is to extend the Bolley–Villani bound to the adapted Wasserstein distance when $\mu$ and $\nu$ are laws of discrete-time processes. This will follow from the adapted version of Theorem 2 which we present in Theorem 7.
@@ 1.1. Main results
The adapted Wasserstein distance $\mathcal{A}\mathcal{W}_{p}$ is well suited for studying optimization problems in which the flow of information plays a central role (see [5]). For example, it induces the smallest topology under which optimal stopping problems have continuous value functions. While the topology induced by $\mathcal{A}\mathcal{W}_{p}$ is well understood, the metric $\mathcal{A}\mathcal{W}_{p}$ remains difficult to control quantitatively. This poses a significant limitation in applications where quantitative control of $\mathcal{A}\mathcal{W}_{p}$ is required.
In this paper, we address this limitation by comparing the adapted weighted total variation with the relative entropy. More precisely, Theorem 7 extends the Bolley–Villani bound (Theorem 2) and shows that the adapted weighted total variation can be bounded in terms of the relative entropy, under suitable moment assumptions. This result provides a new quantitative tool linking an adapted optimal cost to the relative entropy, a fundamental concept in information theory. In particular, Corollary 8 shows that a measure $\mu$ satisfies the adapted $T_{1}$ inequality if and only if $\mu$ has subgaussian tails.
We now introduce the adapted weighted total variation and present our main result, Theorem 7.
@@
where the infimum is attained when $\pi(dx,dy)=(\operatorname{id},\operatorname{id})_{\#}(\mu\wedge\nu)$ on $\{x=y\}$. For example, see [13, Section 2] or [1, Lemma 3.1]. Here, $\mu\wedge\nu$ is the minimum between $\mu$ and $\nu$, defined by
(13) $\mu\wedge\nu(A)=\int_{A}\left(\frac{d\mu}{d(\mu+\nu)}\wedge\frac{d\nu}{d(\mu+\nu)}\right)d(\mu+\nu)$
for all measurable $A$. In this context, $\operatorname{ATV}_{\varphi}$ can be viewed as a natural extension of $\operatorname{TV}_{\varphi}$. The study of $\operatorname{ATV}_{\varphi}$ is not new and related works will be discussed in the next section.
@@
The Bolley–Villani bound in Theorem 2 is a one-step bound that applies at each time step. In the adapted setting, we inductively apply the Bolley–Villani bound to the conditional distributions of $\mu$ to obtain Theorem 7. Thanks to the chain rule of entropy, the Bolley–Villani bound propagates along time without any loss of information. The proofs are provided in Section 2.
By properly choosing the weight function $\varphi$, we obtain the transport-entropy inequality for $\mathcal{A}\mathcal{W}_{1}$. In view of Theorem 1, the adapted $T_{1}$ inequality is equivalent to Gaussian concentration.
@@
Let $\mu\in\mathcal{P}_{1}(\mathcal{X})$ and $C>0$. Each of the following statements implies the next one:
@@
    $\mu$ satisfies the $T_{1}(C)$ inequality.
  3. (c)
    For any $x_{0}\in\mathcal{X}$ and $\alpha\in(0,\frac{1}{2C^{2}})$, $\mu$ satisfies the adapted $T_{1}(C^{\prime})$ inequality where
    (16) $C^{\prime}=\frac{2\sqrt{T}+1}{\sqrt{\alpha}}\left(1+\log\int_{\mathcal{X}}e^{\alpha d(x,x_{0})^{2}}\mu(dx)\right)^{1/2}.$
By the definition of the adapted Wasserstein distance $\mathcal{A}\mathcal{W}_{1}$, it is clear that $\mathcal{W}_{1}\leq\mathcal{A}\mathcal{W}_{1}$ and thus the implication from $(a)$ to $(b)$ follows immediately. The main content of the above corollary lies in the last direction: $(b)$ implies $(c)$.
@@ 1.2. Related works
Let us recall that if the reference measure $\mu$ has independent marginals, i.e., $\mu=\otimes_{t=1}^{T}\gamma_{t}$ and each $\gamma_{t}$ satisfies the $T_{1}$ inequality (2) with a constant $C>0$, then the tensorization argument yields
@@
Our result is based on the Bolley–Villani bound (Theorem 2) and does not extend the tensorization technique beyond the independent case. We aim to investigate its generalization in future research.
@@
(24) $\operatorname{TV}(\mu,\nu):=2\inf_{\pi\in\operatorname{Cpl}(\mu,\nu)}\pi(x\neq y)\leq\sqrt{2H\left(\nu\,|\,\mu\right)}.$
@@
holds with $C=2^{T}-1$. The optimal constant C=2T1C=2T-1 was later obtained by Acciaio et al. [1]. In combination with the classical Pinsker inequality, this implies
@@
Acciaio et al. [1] further showed that if $\mu\in\mathcal{P}((\mathbb{R}^{d})^{T})$ satisfies
(30) $\mu\text{-ess sup}\frac{1}{1+\sum_{s=1}^{t}\left|x_{s}\right|}\int\left|x_{t+1}\right|\mu^{x_{1},\ldots,x_{t}}(dx_{t+1})<\infty\text{ for all }t\in\{1,2,\ldots,T-1\},$
@@
for some constant $C>0$ that arises from the assumption (30). This estimate, when combined with the Bolley–Villani bound (Theorem 2), naturally leads to the following transport inequality:
@@ 2. Proofs
If $\gamma$ is a finite signed measure, we write its Jordan decomposition as $\gamma=\gamma_{+}-\gamma_{-}$, where $\gamma_{+}$ and $\gamma_{-}$ are the positive and negative variations of $\gamma$, respectively.
@@
where $\eta_{1}:=\mu_{1}\wedge\nu_{1}$. Recall that $(\mu_{1}-\nu_{1})_{+}$ and $(\nu_{1}-\mu_{1})_{+}$ denote the positive variations of $\mu_{1}-\nu_{1}$ and $\nu_{1}-\mu_{1}$, respectively. For each $t\in\{1,2,\ldots,T-1\}$ and $x_{1:t},y_{1:t}\in\mathcal{X}_{1:t}$, we define $\gamma^{x_{1:t},y_{1:t}}\in\operatorname{Cpl}(\mu^{x_{1:t}},\nu^{y_{1:t}})$ via
@@
The second inequality follows from the minimality of $\mu_{1:j-1}\wedge\nu_{1:j-1}$. Indeed,
(63) $\eta^{x_{1:j-2}}(dx_{j-1})\cdots\eta^{x_{1}}(dx_{2})\eta_{1}(dx_{1})\leq\mu^{x_{1:j-1}}(dx_{j-1})\cdots\mu^{x_{1}}(dx_{2})\mu_{1}(dx_{1})=\mu_{1:j-1}(dx_{1:j-1})$
and similarly,
(64) $\eta^{x_{1:j-2}}(dx_{j-1})\cdots\eta^{x_{1}}(dx_{2})\eta_{1}(dx_{1})\leq\nu_{1:j-1}(dx_{1:j-1}).$
Hence,
(65) $\eta^{x_{1:j-2}}(dx_{j-1})\cdots\eta^{x_{1}}(dx_{2})\eta_{1}(dx_{1})\leq\mu_{1:j-1}\wedge\nu_{1:j-1}(dx_{1:j-1}).$
The proof of the last case $j=T$ is essentially the same. Indeed,
@@
(67) $\leq\int_{\mathcal{X}}\varphi(x)\left|\mu^{x_{1:T-1}}-\nu^{x_{1:T-1}}\right|(dx_{T})\mu_{1:T-1}\wedge\nu_{1:T-1}(dx_{1:T-1})$
@@
(84) $\leq\left(\int_{\mathcal{X}_{1:j-1}}f(x_{1:j-1})^{1/2}g(x_{1:j-1})^{1/2}\mu_{1:j-1}\wedge\nu_{1:j-1}(dx_{1:j-1})\right)^{2}$
@@
By applying Theorem 7 to $\varphi(x):=\sqrt{\alpha}d(x,x_{0})^{p}$, we immediately establish the estimates stated in Corollary 8 and Corollary 9. In particular, if $\int_{\mathcal{X}}e^{\alpha d(x,x_{0})^{2}}\mu(dx)<\infty$ for some $x_{0}\in\mathcal{X}$ and $\alpha>0$, then $\mu$ satisfies the adapted $T_{1}$ inequality. This proves the implication from $(b)$ to $(c)$ in Corollary 8. The other direction is immediate from $\mathcal{W}_{1}\leq\mathcal{A}\mathcal{W}_{1}$.